If at least 1 child in a family with 3 children is a boy find the prob that family have 2 of the children are boys


1) 3/7


2) 4/7


3) 3/8


4) 1/3

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

Given at least 1 child in a family with 3 children is a boy

so event  1 boy in a family  ,2 boy in a family ,  3 boy in a fimily are equally likily event

 so chances of happening any event fou of three is 1/3

 so prob that family have 2 of the children are boys =1/3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

swetha viriyala
13 years ago

There is a probability that the family contains 1 boy,2boys or 3 boys.... so chance of each event selection is 1/3 but its not the final answer..

if one boy is there then chance of selecting that 1 boy from 3 children is 1/3 

if 2 boys are there then chance of selecting one boy among them is 2/3

if 3 boys are there then chance of selecting 1 boy is 3/3 i.e., 1

Now, required probability is 1/3 * 1 / 1/3 * 1 + 1/3 * 2/3 +1/3 * 1/3 = 1/2 

This is the final answer required.